Factorize (X^10 + Y^10) Please solve in detailed way. Thanks.

Factorize
 
(X^10 + Y^10)
 
 
 
Please solve in detailed way. Thanks. 
 

Grade:9

3 Answers

Arun
4 years ago
Dear Anuj
 
I think your question is wrong. Please check and post the correct question. You can also attach an image. I will be happy to help you.
Vikas TU
4 years ago
Dear student 
Question should be , Factorise (X^10 –  Y^10)
It can be written as :
(x5)2 – (y10)
Rewrite same with y .
(x5)2 – (y5)2
Expand it by using formula (a+b)(a-b)
(x5 + y5 )(x5 –  y5 )
Aditya Gupta
4 years ago
actually x^10+y^10 can be factorised.
write it as (x^2)^5 + (y^2)5
but we know the identity a^(2n+1) + b^(2n+1)= (a+b)[a^2nb^0 – a^(2n-1)b^1+....... – a^1b^(2n-1)+a^0b^2n]
in this identity plug a= x^2, b= y^2 and n=2 and you shall obtain the reqd factorisation.
kindly approve :)